THEORETICAL CALCULATION OF THE N_{2} BROADENED HALF-WIDTHS OF H_{2}O

Q. MA, NASA/Goddard Institute for Space Studies and Department of Applied Physics and Applied methamatics, Columbia University, 2880 Broadway, New York, NY 10025; \underline{R. H. TIPPING}, Department of Physics and Astronomy, University of Alabama, Tuscaloosa, AL 35487; R. R. GAMACHE, Department of Environmental, Earth and Atmospheric Science, University of Mass. Lowell, Lowell, MA 01854.

The water molecule is the most important Greenhouse gas and thus plays a pivotal role in atmospheric spectra. In addition to accurate intensities and frequencies, one also needs accurate self and foreign half-widths and shifts, and their temperature dependence. Over the years, a large number of theoretical calculations have been carried out by Gamache and his collaborators^{a}. They used the complex Robert-Bonamy theory with a sophisticated interaction potential. The drawback of this method is that one has to carry out the calculations to a high-order perturbation in order to obtain converged results. However, by using the coordinate representation one is able to obviate the perturbation expansion and obtain results corresponding to a high cut-off order^{b}. We present comparisons for the H_{2}O-N_{2} system for a few lines using the same interaction potential for a comparison between the methods. We conclude that for lines having a large half-width, the convergence is rapid but, on the other hand, for lines with relatively small half-widths the convergence is very slow.

aI. E. Gordon et al., J.Q.S.R.T. 108, 389 (2007) and references therein.

bQ. Ma, R. H. Tipping, and C. Boulet, J. Chem. Phys. 124, 014109 (2006).